OpenStudy (anonymous):

If 5 apples in a barrel of 25 apples are rotten, what is the expected number of rotten apples in a sample of 2 apples?

OpenStudy (anonymous):

\[2\times \frac{5}{25}\]

OpenStudy (anonymous):

A. 0.4 B. 1 C. 0.63 D. 0.33

OpenStudy (anonymous):

another straight forward one, since the probabilty of getting a rotton one is \(\frac{5}{25}=\frac{1}{5}\)

OpenStudy (anonymous):

multiply that by the number of apples you choose, namely 2 and get \(\frac{2}{5}\)

OpenStudy (anonymous):

Decimal ? (: